Circular No. 6156

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

X-RAY SOURCES IN M31 G. L. Israel and A. Treves, International School for Advanced Studies

(SISSA-ISAS), Trieste; L. Stella and G. Tagliaferri, Osservatorio Astronomico di Brera, Merate; and T. Belloni, Anton Pannekoek Astronomical Institute, Amsterdam, communicate: “We performed a power-spectrum analysis of the archival ROSAT High Resolution Imager (HRI) light curves of the thirteen brightest x-ray sources detected in M31 during a long (48000s) exposure on 1990 July 25–28 (cf. Primini et al. 1993, Ap.J. 410, 615). Coherent pulsations at a barycentric period of 76.92(3) s are detected in ∼ 600 photons from ROSAT source 58 (= Einstein source 69; cf. Trinchieri and Fabbiano 1991, Ap.J. 382, 82) with a significance > 3.4σ; the modulation is nearly sinusoidal in shape, with a 40-percent pulsed fraction (i.e., semiamplitude modulation divided by the mean source count rates). We find also some evidence for 3.01856(5)-s periodic modulation in ∼ 300 photons from ROSAT source 60 (Einstein source 71), with a 50-percent pulsed fraction and a significance > 2.7σ. We also searched for periodicities in PSPC data of the two sources obtained on 1991 July 27 (30000 s), 1992 Aug. 6–9 (10000 s), and 1993 Jan. 2–7 (17000 s). No significant pulsations were found; the corresponding upper limits to the amplitude were in all cases consistent with the HRI results. Both sources have luminosities close to 10^{38} erg/s, such that large period derivatives are expected. If confirmed, these would be the first x-ray pulsators detected beyond the Magellanic Clouds. Follow-up x-ray and optical observations are encouraged.”

SUPERNOVA 1995D IN NGC 2962 R. W. Argyle and L. V. Morrison, Royal Greenwich Observatory, report

an accurate optical position for SN 1995D obtained with the Carlsberg Automatic Meridian Circle on La Palma, operated by P. S. Herridge and F. Belizon: α = 9^{h}40^{m}54^{s}.753 ± 0^{s}.005, δ = +5^{o}08′26′′.16 ± 0′′.08 from 8 observations (equinox J2000.0, FK5 reference frame, epoch 1995.16). The following V magnitudes have also been obtained (± 0.13 mag): Feb. 23.03 UT, 13.46; 24.03, 13.42; 25.03, 13.39; 26.02, 13.54.

Visual magnitude estimates (cf. IAUC 6137): Mar. 21.793 UT, 14.2 (L. Szentaskó, Veresegyház, Hungary); 24.82, 14.6 (K. Sárneczky, Ráktanya, Hungary); 28.83, 14.6 (Szentaskó); Apr. 2.79, 15.0 (Sárneczky).

1995 April 6 Daniel W. E. Green